OpenStudy (anonymous):

I have to solve the system; y = 1/2x - 6 2x + 6y = 19 The fact they are in different forms is throwing me off.. help?!

OpenStudy (anonymous):

You can find x first, bc y is given

OpenStudy (anonymous):

So just like substitution?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

\[2x+6\left(\frac{1}{2}x-6\right)=19\]

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

Do you know how I can solve this in a nother way?